# Projecting Gaussian Ellipsoids While Avoiding Affine Projection Approximation

### v1

@@ Abstract
Recently, 3D Gaussian Splatting has dominated novel-view synthesis with its real-time rendering speed and state-of-the-art rendering quality. However, during the rendering process, the use of the Jacobian of the affine approximation of the projection transformation leads to inevitable errors, resulting in blurriness, artifacts and a lack of scene consistency in the final rendered images. To address this issue, we introduce an ellipsoid-based projection method to calculate the projection of Gaussian ellipsoid on the image plane, witch is the primitive of 3D Gaussian Splatting. As our proposed ellipsoid-based projection method cannot handle Gaussian ellipsoids with camera origins inside them or parts lying below $z=0$ plane in the camera space, we designed a pre-filtering strategy. Experiments over multiple widely adopted benchmark datasets show that using our ellipsoid-based projection method can enhance the rendering quality of 3D Gaussian Splatting and its extensions.
Figure 1: Our method achieves a comprehensive improvement in rendering quality compared to 3D Gaussian splatting (3D-GS) [24] while also accelerating rendering speed. We propose an ellipsoid-based projection method to replace the Jacobian of the affine approximation of the projection transformation in 3DGS. Our ellipsoid-based projection method can be applied to any 3DGS-based work to enhance rendering quality. This figure shows the rendering results of applying our method to 3DGS and Mip-Splatting [52] in the scene bicycle of Mip-NeRF360 dataset [3], in witch the rendering quality is enhanced with much less blur and artifacts.
@@ 1 Introduction
Novel View Synthesis (NVS) plays a crucial role in computer vision and computer graphics, with numerous applications, including robotics, virtual reality, and 3D gaming. One of the most influential works in this field is the Neural Radiance Field (NeRF) [36], proposed by Mildenhall et al. in 2020. NeRF utilizes a Multilayer Perceptron (MLP) to store geometric and appearance information of a scene and employs differentiable volume rendering [9, 29, 35, 22]. Although NeRF and its extensions can render high-quality images, their training time is excessively long, and the rendering speed is far from meeting the standards of real-time rendering ($\geq$ 30 fps). Recently, 3D Gaussian Splatting (3DGS) [24] has made a significant impact in NVS due to its real-time rendering speed, high-quality rendering results and competitive training times. Unlike NeRF, 3DGS represents scenes using a set of Gaussian ellipsoids explicitly. By projecting each Gaussian ellipsoid onto the image plane and using $\alpha$-blending for rendering, the properties of each Gaussian ellipsoid, position, pose, scale, transparency, and color, are optimized based on a multi-view photometric loss.
Although 3D Gaussian Splatting has shown impressive results, due to the local affine approximation of the projection transformation at the center of each Gaussian ellipsoid during rendering, errors are inevitably introduced, negatively affecting rendering quality and scene consistency. We observed blurriness and artifacts in distant objects in the scene, which we attribute to larger errors in the approximated projection transformation as the distance from the Gaussian ellipsoid center increases, particularly since distant Gaussian ellipsoids are generally large. Additionally, while the details rendered in the training set are of high quality, the results in the test set show a decline, likely due to the lack of scene consistency on account of the approximated projection transformation.
@@
During our experiments, we discovered that there are two types of Gaussian ellipsoids that will cause the training process to diverge and negatively impact the system. The first type has the camera’s origin inside the ellipsoid, where no lines through the camera’s origin can be tangent to it. The second type consists of Gaussian ellipsoids that have a portion below $z=0$ plane in the camera space, the projection of these ellipsoids results in hyperbola or parabola [10] rather than an ellipse. To avoid negatively impacting the system, we designed filtering algorithms specifically for these two types of Gaussian ellipsoids. Extensive experiments demonstrated that our method not only improves rendering quality compared to 3DGS but also further accelerates rendering speed.
@@
  • We proposed an ellipsoid-based projection method, addressing the negative impact on rendering quality and scene consistency caused by approximating the projection transformation using the Jacobian of its affine approximation in 3DGS.
@@
  • Experiments conducted on challenging benchmark datasets demonstrated that our method surpasses 3DGS in both rendering quality and rendering speed.
  • We only modified the code related to the projection part in 3DGS, making our ellipsoid-based projection method easily transferable to other works based on 3DGS to further improve their rendering quality.
@@ 2.1 Novel View Synthesis
The goal of Novel View Synthesis (NVS) is to generate images from new perspectives different from those of the captured images. There have been notable progressions in NVS, especially since the introduction of Neural Radiance Fields (NeRF) [36]. NeRF uses a Multi-Layer Perceptron (MLP) to represent represent geometry and view-dependent appearance, optimized through volume rendering techniques [9, 29, 35, 22] to achieve high-quality rendering results. However, during the rendering process, the geometric and appearance information for each point along the ray must be obtained through a complex MLP network, resulting in slow rendering speeds. Subsequent work has utilized distillation and baking techniques [20, 41, 42, 49, 51] to improve NeRF’s rendering speed, usually sacrificing rendering quality. Furthermore, using feature-grid representations [38, 6, 14, 45, 26] have increased the training speed of NeRF. In addition, some extended works have further improved rendering quality [2, 3, 4] achieving state-of-the-art performance.
@@ 2.2 Primitive-based Differentiable Rendering
Primitive-based rendering techniques, which project geometric primitives onto the image plane, have been widely studied for their efficiency [16, 17, 39, 43, 56, 57]. Differentiable point-based rendering methods provide significant flexibility in representing complex structures, making them ideal for novel view synthesis. Notably, NPBG [1] rasterizes features from point clouds onto the image plane and then employs a convolutional neural network for RGB image prediction. DSS [50] aims to optimize oriented point clouds derived from multi-view images under specific lighting conditions. Pulsar [27] presents a tile-based acceleration structure to enhance the efficiency of rasterization. More recently, 3DGS [24] projects anisotropic Gaussian ellipsoids onto the image plane and uses $\alpha$-blending techniques to render the ellipses on the plane based on depth ordering, to further optimizing the Gaussian ellipsoids. During the projection process, 3DGS introduces errors by approximating the projection transformation using the Jacobian of its affine approximation. To fundamentally eliminate the negative impact of these errors, we propose an ellipsoid-based projection method to replace the Jacobian of the affine approximation of the projection transformation in 3DGS.
@@ 2.3 Perspective Projection of an Ellipsoid
In 1999, David Eberly et al. provided a comprehensive process for calculating the projection of an ellipsoid onto a given plane [10]. This algorithm has primarily been used for the localization of 3D objects [44, 11, 15] and for designing shadow models [5, 32]. In this work, we build upon David Eberly’s findings to propose a projection method for projecting 3D Gaussian ellipsoids onto the image plane. The representation of the ellipsoid is achieved using the inverse of the covariance matrix of 3D gaussian function. After obtaining the ellipse equation, we further convert the result from the camera space to the image space. The projection algorithm [10] proposed by David Eberly et al. has certain preconditions. Therefore, we designed a pre-filtering strategy to eliminate Gaussian ellipsoids that do not meet these preconditions before the rendering process.
@@ 3 Preliminaries
3D Gaussian splatting [24] proposes a method for representing three-dimensional scenes using Gaussian ellipsoids as scene primitives. The geometric features of each Gaussian ellipsoid are represented by $\mathbf{p}\in\mathbb{R}^{3\times 1}$ and the covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$. Where, $\mathbf{p}$ controls the geometric position of the Gaussian ellipsoid, while $\mathbf{\Sigma}$ controls its shape. The covariance matrix $\mathbf{\Sigma}$ consist of a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$, which control the orientation and scale of the Gaussian ellipsoid, respectively.
@@
3DGS generates a set of sparse point clouds through Structure from Motion (SfM) to initialize the Gaussian ellipsoids, with the centers of the ellipsoids corresponding to the point clouds. For further rendering, the covariance matrix $\mathbf{\Sigma}$ of a Gaussian ellipsoid is transformed into the camera space using a viewing transformation $\mathbf{W}$. Then, the covariance matrix is projected to image plane via the projection transformation, which is approximated by the corresponding local affine approximation Jacobian matrix $\mathbf{J}$.
@@
After removing the third row and third column of $\mathbf{\Sigma}^{\prime}$, the 2D covariance matrix $\mathbf{\Sigma}^{2D}$ was obtained. By projecting the center point of the Gaussian ellipsoid to the image plane using the viewing transformation $\mathbf{W}$ and the projection transformation $\mathbf{P}$, the center point of the ellipse $\mathbf{p}^{2D}$ can be obtained.
$\mathbf{p}^{2D}=\mathbf{PWp}$ (3)
@@ 3.2 Projection Algorithm for Ellipsoid
The standard equation of an ellipsoid is $(\mathbf{x}-\mathbf{c})^{T}\mathbf{A}(\mathbf{x}-\mathbf{c})=1$, where $\mathbf{c}$ is the center of the ellipsoid and $\mathbf{A}$ is a positive definite matrix. Given the camera origin $\mathbf{e}$ and the image plane $\mathbf{n}\cdot\mathbf{x}=d$, where $\mathbf{n}$ is a unit-length vector, $d$ is a constant. The task is to calculate the projection of the ellipsoid onto the image plane.
@@
The direction vector $D$ of the ray can also be expressed as $\mathbf{d}=\frac{\mathbf{x}-\mathbf{k}}{|\mathbf{x}-\mathbf{k}|}$. Substituting this into the equation yields
$(\mathbf{x}-\mathbf{k})^{T}(\mathbf{A}^{T}\mathbf{\delta}\mathbf{\delta}^{T}\mathbf{A}-(\mathbf{\delta}^{T}\mathbf{A}\mathbf{\delta}-1)\mathbf{A})(\mathbf{x}-\mathbf{k})=0.$ (10)
This equation describes a cone with vertex at E.
Next, by solving the combined equations of the cone and the image plane, the projection of the ellipsoid onto the plane can be obtained.
@@ 4 Proposed Method
Due to errors introduced by approximating the projection transformation using the Jacobian of its affine approximation in 3D Gaussian splatting [24], we propose an ellipsoid-based projection method to eliminate the errors and enhance rendering quality and scene consistency. In Sec. 4.1, we explain why errors are inevitably introduced in the projection process of 3DGS, followed by the presentation of our ellipsoid-based projection method and its theoretical derivation. Next, in Sec. 4.2, we identify two types of Gaussian ellipsoids that cannot be rendered using our proposed projection method and present a pre-filtering strategy to filter out these two types of Gaussian ellipsoids.
@@ 4.1 Ellipsoid-based Projection Method
Gaussian functions are closed under linear transformations but not under nonlinear transformations, which means the result of nonlinear transformations can no longer represent an ellipsoid (or ellipse). Since projection transformation is nonlinear, 3DGS [24] adopts a local affine approximation of the projection transformation $\mathbf{P}$ at the center of the Gaussian ellipsoid to obtain the Jacobian matrix $\mathbf{J}$. Using the the Jacobian of the affine approximation of the projection transformation will inevitably introduce errors at positions other than the center of the Gaussian ellipsoid during the projection process, with errors increasing the farther one moves from the center.
@@
Based on the method introduced by David Eberly et al. [10], we designed an ellipsoid-based projection method for 3DGS. To simplify the computation, we will project the Gaussian ellipsoid in the camera space $\mathcal{C}\{x,y,z\}$. First, we need to derive the equation of the Gaussian ellipsoid based on the 3D covariance matrix $\mathbf{\Sigma}$. In 3D Gaussian Splatting [24], the author obtains the corresponding Gaussian ellipse equation by setting the exponent part of the 2D Gaussian function equal to $3^{2}$ (satisfying the $3\sigma$ principle). Similarly, for the 3D Gaussian function, we also set its exponent part equal to $3^{2}$:
@@
where $\mathbf{p}_{c}$ is the center point of Gaussian ellipsoid in camera space and $\mathbf{\Sigma}_{c}$ is the 3D covariance matrix $\mathbf{\Sigma}$ in camera space.
In camera space, the camera origin $\mathbf{c}=[0,0,0]$. According to equations Eq. 6, Eq. 8 and Eq. 10, the equation of the cone can be obtained.
@@
Setting the third element of $\mathbf{x}=[x,y,z]$ to 1 gives the intersection between the cone and the $z=1$ plane, and the equation of this intersection is an ellipse equation that depends only on $x$ and $y$.
For a point $\mathbf{x}^{2d}=[x,y]$ on the $z=1$ plane, scaling and translation are required to convert it into a point $\mathbf{x}^{img}=[x^{img},y^{img}]$ on the image plane $\mathcal{I}\{x^{img},y^{img}\}$. The corresponding relationship is
@@
in witch $f_{x}$ and $f_{y}$ are the camera intrinsic parameters, and $w$ and $h$ represent the width and height of the image, respectively.
@@
where $\mathbf{\Sigma}_{img}^{-1}\in\mathbb{R}^{2\times 2}$ is the inverse of the second-order covariance matrix corresponding to the ellipse and $\mathbf{p}_{img}\in\mathbb{R}^{2\times 1}$ is the center point of the ellipse.
We find that the center point of the ellipse calculated from our projection transformation has a slight difference compared to the result obtained by directly projecting the center point of the Gaussian ellipsoid onto the image plane. This indicates that the center point of the ellipse is no longer the projection of the center point of the Gaussian ellipsoid, which also explains why the projected result remains an ellipse, but it is not possible to find an affine transformation that directly converts the 3D Gaussian covariance matrix into a 2D one.

### v3

@@ Abstract
Recently, 3D Gaussian Splatting has dominated novel-view synthesis with its real-time rendering speed and state-of-the-art rendering quality. However, during the rendering process, the use of the Jacobian of the affine approximation of the projection transformation leads to inevitable errors, resulting in blurriness, artifacts and a lack of scene consistency in the final rendered images. To address this issue, we introduce an ellipsoid-based projection method to calculate the projection of Gaussian ellipsoid onto the image plane, which is the primitive of 3D Gaussian Splatting. As our proposed ellipsoid-based projection method cannot handle Gaussian ellipsoids with camera origins inside them or parts lying below $z=0$ plane in the camera space, we designed a pre-filtering strategy. Experiments over multiple widely adopted benchmark datasets show that our ellipsoid-based projection method can enhance the rendering quality of 3D Gaussian Splatting and its extensions.
Figure 1: Our method achieves a comprehensive improvement in rendering quality and rendering speed compared to 3D Gaussian splatting (3DGS) [24]. We propose an ellipsoid-based projection method to replace the Jacobian of the affine approximation of the projection transformation in 3DGS. Our ellipsoid-based projection method can be applied to any 3DGS-based work to enhance rendering quality. This figure shows the rendering results of applying our method to 3DGS and Mip-Splatting [52] in the scene bicycle of Mip-NeRF360 dataset [3], in which the rendering quality is enhanced with less blur and artifacts.
@@ 1 Introduction
Novel View Synthesis (NVS) plays a crucial role in computer vision and computer graphics, with numerous applications, including robotics, virtual reality, and 3D gaming. One of the most influential works in this field is the Neural Radiance Field (NeRF) [36], proposed by Mildenhall et al. in 2020. NeRF utilizes a Multilayer Perceptron (MLP) to store geometric and appearance information of a scene and employs differentiable volume rendering [9, 29, 35, 22]. Although NeRF and its extensions can render high-quality images, their training time is excessively long, and the rendering speed is far from achieving real-time rendering ($\geq$ 30 fps). Recently, 3D Gaussian Splatting (3DGS) [24] has made a significant impact in NVS due to its real-time rendering speed, high-quality rendering results and competitive training times. Unlike NeRF, 3DGS represents scenes with a set of Gaussian ellipsoids explicitly. By projecting each Gaussian ellipsoid onto the image plane and using $\alpha$-blending for rendering, the properties of each Gaussian ellipsoid, position, pose, scale, transparency, and color, are optimized based on a multi-view photometric loss.
Although 3D Gaussian Splatting has shown impressive results, due to the local affine approximation of the projection transformation [58] at the center of each Gaussian ellipsoid during rendering, errors are inevitably introduced, negatively affecting rendering quality and scene consistency. We observed blurriness and artifacts in distant objects in the scene, which we attribute to larger errors in the approximated projection transformation as the distance from the Gaussian ellipsoid center grows, particularly when distant Gaussian ellipsoids are generally large. Additionally, while the details rendered in the training set are of high quality, the results in the test set show a decline, likely due to the lack of scene consistency on account of the approximated projection transformation.
@@
The following experiments demonstrate that there are two types of Gaussian ellipsoids that can cause the training process to diverge and negatively impact the system. The first type has the camera’s origin inside the ellipsoid, where no lines through the camera’s origin can be tangent to it. The second type consists of Gaussian ellipsoids that have a portion below $z=0$ plane in the camera space, the projection of these ellipsoids results in hyperbola or parabola [10] rather than an ellipse. To avoid negatively impacting the system, we designed filtering algorithms specifically for these two types of Gaussian ellipsoids. Extensive experiments demonstrated that our method not only improves rendering quality compared to 3DGS but also further accelerates rendering speed.
@@
  • We proposed an ellipsoid-based projection method to eliminate the negative impact on rendering quality and scene consistency caused by approximating the projection transformation using the Jacobian of its affine approximation in 3DGS.
@@
  • Experiments conducted on challenging benchmark datasets demonstrated that our method surpasses 3DGS in both rendering quality and speed.
  • Our ellipsoid-based projection method shows improvement results on 3DGS and its extensions, and can be easily applied to them, requiring only few changes to the original code.
@@ 2.1 Novel View Synthesis
The goal of Novel View Synthesis (NVS) is to generate images from new perspectives different from those of the captured images. There have been notable progressions in NVS, especially since the introduction of Neural Radiance Fields (NeRF) [36]. NeRF uses a Multi-Layer Perceptron (MLP) to represent geometry and view-dependent appearance, optimized through volume rendering techniques [9, 29, 35, 22] to achieve high-quality rendering results. However, during the rendering process, the geometric and appearance information for each point along the ray must be obtained through a complex MLP network, resulting in slow rendering speed. Subsequent work has utilized distillation and baking techniques [20, 41, 42, 49, 51] to increase NeRF’s rendering speed, usually at a cost of lower rendering quality. Furthermore, the training speed and representational power of NeRF have been enhanced using feature-grid based scene representations [38, 6, 14, 45, 26]. In addition, some extended works have further improved rendering quality [2, 3, 4] achieving state-of-the-art performance.
@@ 2.2 Primitive-based Differentiable Rendering
Primitive-based rendering techniques, which project geometric primitives onto the image plane, have been widely studied for their efficiency [16, 17, 39, 43, 56, 57]. Differentiable point-based rendering methods provide significant flexibility in representing complex structures, making them ideal for novel view synthesis. Notably, NPBG [1] rasterizes features from point clouds onto the image plane and then employs a convolutional neural network for RGB image prediction. DSS [50] aims to optimize oriented point clouds derived from multi-view images under specific lighting conditions. Pulsar [27] presents a tile-based acceleration structure to enhance the efficiency of rasterization. More recently, 3DGS [24] projects anisotropic Gaussian ellipsoids onto the image plane and uses $\alpha$-blending techniques to render the ellipses on the plane based on depth ordering, to further optimize the Gaussian ellipsoids. During the projection process, 3DGS introduces errors by approximating the projection transformation using the Jacobian of its affine approximation [58]. To fundamentally eliminate the negative impact of these errors, we propose an ellipsoid-based projection method to replace the Jacobian of the affine approximation of the projection transformation in 3DGS.
@@ 2.3 Perspective Projection of an Ellipsoid
In 1999, David Eberly et al. provided a comprehensive process for calculating the projection of an ellipsoid onto a given plane [10]. This algorithm has primarily been used for the localization of 3D objects [44, 11, 15] and for designing shadow models [5, 32]. In this work, based on David Eberly’s findings, we propose a projection method for projecting 3D Gaussian ellipsoids onto the image plane. The representation of the ellipsoid is achieved by using the inverse of the covariance matrix of 3D gaussian function. After obtaining the ellipse equation, we further convert the result from the camera space to the image space. The projection algorithm [10] proposed by David Eberly et al. has certain preconditions to ensure the projection of the ellipsoid onto the image plane is an ellipse. Therefore, we designed a pre-filtering strategy to eliminate Gaussian ellipsoids that do not meet these preconditions before the rendering process.
@@ 3 Preliminaries
3D Gaussian splatting [24] represents three-dimensional scenes using Gaussian ellipsoids as scene primitives. The geometric features of each Gaussian ellipsoid are represented by $\mathbf{p}\in\mathbb{R}^{3\times 1}$ and the covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$. Where, $\mathbf{p}$ controls the geometric position of the Gaussian ellipsoid, and $\mathbf{\Sigma}$ controls its shape. The covariance matrix $\mathbf{\Sigma}$ consist of a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$, which controls the orientation and the scale of the Gaussian ellipsoid, respectively.
@@
3DGS generates a set of sparse point clouds through Structure from Motion (SfM) to initialize the Gaussian ellipsoids, with centers of the ellipsoids corresponding to the point clouds. For further rendering, the covariance matrix $\mathbf{\Sigma}$ of a Gaussian ellipsoid is transformed into the camera space using a viewing transformation $\mathbf{W}$. Then, the covariance matrix is projected onto the image plane via the projection transformation, which is approximated by the corresponding local affine approximation Jacobian matrix $\mathbf{J}$.
@@
After removing the third row and third column of $\mathbf{\Sigma}^{\prime}$, the 2D covariance matrix $\mathbf{\Sigma}^{2D}$ was obtained. By projecting the center of the Gaussian ellipsoid onto the image plane using the viewing transformation $\mathbf{W}$ and the projection transformation $\mathbf{P}$, the center of the ellipse $\mathbf{p}^{2D}$ can be obtained.
@@ 3.2 Projection Algorithm for Ellipsoid
The standard equation of an ellipsoid is $(\mathbf{x}-\mathbf{c})^{T}\mathbf{A}(\mathbf{x}-\mathbf{c})=1$, where $\mathbf{c}$ is the center of the ellipsoid and $\mathbf{A}$ is a positive definite matrix. Given the camera origin $\mathbf{e}$ and the image plane $\mathbf{n}\cdot\mathbf{x}=k$, where $\mathbf{n}$ is a unit-length vector, $k$ is a constant. The task is to calculate the projection of the ellipsoid onto the image plane.
@@
The direction vector $\mathbf{d}$ of the ray can also be expressed as $\mathbf{d}=\frac{\mathbf{x}-\mathbf{e}}{|\mathbf{x}-\mathbf{e}|}$. Substituting this into the equation yields
$(\mathbf{x}-\mathbf{e})^{T}(\mathbf{A}^{T}\mathbf{\delta}\mathbf{\delta}^{T}\mathbf{A}-(\mathbf{\delta}^{T}\mathbf{A}\mathbf{\delta}-1)\mathbf{A})(\mathbf{x}-\mathbf{e})=0.$ (9)
This equation describes a cone with vertex at $\mathbf{e}$.
Next, by solving the combined equations of the cone and the image plane, the projection of the ellipsoid onto the image plane can be obtained.
@@ 4 Proposed Method
Due to errors introduced by approximating the projection transformation using the Jacobian of its affine approximation in 3D Gaussian splatting [24], we propose an ellipsoid-based projection method to eliminate the errors and enhance rendering quality and scene consistency. In Sec. 4.1, we explain why errors are inevitably introduced in the projection process of 3DGS, followed by the presentation of our ellipsoid-based projection method and its theoretical derivation. Next, in Sec. 4.2, we identify two types of Gaussian ellipsoids that cannot be rendered with the proposed projection method and present a pre-filtering strategy to filter out these two types of Gaussian ellipsoids.
@@ 4.1 Ellipsoid-based Projection Method
Gaussian functions are closed under linear transformations but not under nonlinear transformations, which means the result of nonlinear transformations no longer represent an ellipsoid (or ellipse). Since projection transformation is nonlinear, 3DGS [24] adopts a local affine approximation of the projection transformation $\mathbf{P}$ at the center of the Gaussian ellipsoid to obtain the Jacobian matrix $\mathbf{J}$. Using the Jacobian of the affine approximation of the projection transformation will inevitably introduce errors at positions other than the center of the Gaussian ellipsoid during the projection process, and farther distance from the center gives larger error.
@@
Based on the method introduced by David Eberly et al. [10], we designed an ellipsoid-based projection method for 3DGS. To simplify the computation, we will project the Gaussian ellipsoid in the camera space $\mathcal{C}\{x,y,z\}$. First, we need to derive the equation of the Gaussian ellipsoid based on 3D covariance matrix $\mathbf{\Sigma}$. In 3D Gaussian Splatting [24], the author obtains the corresponding Gaussian ellipse equation by setting the exponential part of the 2D Gaussian function equal to $3^{2}$ (satisfying the $3\sigma$ principle). Similarly, for the 3D Gaussian function, we also set its exponential part equal to $3^{2}$:
@@
where $\mathbf{p}_{c}$ is the center of Gaussian ellipsoid in camera space and $\mathbf{\Sigma}_{c}$ is the 3D covariance matrix $\mathbf{\Sigma}$ in camera space.
In camera space, the camera origin $\mathbf{c}=[0,0,0]$. According to Eq. 5, Eq. 7 and Eq. 9, the equation of the cone can be obtained.
@@
Setting the third element of $\mathbf{x}=[x,y,z]$ to 1 gives the intersection between the cone and the $z=1$ plane. Thus, the equation of this intersection is an ellipse equation that depends only on $x$ and $y$.
For one point $\mathbf{x}^{2d}=[x,y]$ in the $z=1$ plane, scaling and translation are required to convert it into a point $\mathbf{x}^{img}=[x^{img},y^{img}]$ in the image plane $\mathcal{I}\{x^{img},y^{img}\}$. The corresponding relationship is
@@
where $f_{x}$ and $f_{y}$ are the camera intrinsic parameters, and $w$ and $h$ represent the width and height of the image, respectively.
@@
where $\mathbf{\Sigma}_{img}^{-1}\in\mathbb{R}^{2\times 2}$ is the inverse of the second-order covariance matrix corresponding to the ellipse and $\mathbf{p}_{img}\in\mathbb{R}^{2\times 1}$ is the center of the ellipse.
We find that the center of the ellipse calculated from our projection transformation has a slight difference compared to the result obtained by directly projecting the center of the Gaussian ellipsoid onto the image plane. This indicates that the center of the ellipse is no longer the projection of the center of the Gaussian ellipsoid, which also explains why the projected result remains an ellipse, but it is not possible to find an affine transformation that directly converts the 3D Gaussian covariance matrix into a 2D one.